Circular No. 7509

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000du IN UGC 3920 Y. Sato and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Oct. 14.5 (mag ∼ 17.8) and 16.5 UT (mag ∼ 17.7). The new object is located at α = 7^{h}35^{m}19^{s}.78, δ = +19^{o}02′43′′.5 (equinox 2000.0), which is 14′′.5 east and 1′′.5 south of the nucleus of UGC 3920. A KAIT image of the same field taken on Apr. 5.1 showed nothing at the position of SN 2000du (limiting mag ∼ 19.0).

SUPERNOVA 2000dv IN UGC 4671 Li also reports the discovery of an apparent supernova by M. Pa-

penkova on unfiltered KAIT images taken on Oct. 17.5 and 18.5 UT, both showing the new object at mag ∼ 18.0. SN 2000dv is located at α = 8^{h}56^{m}43^{s}.40, δ = +52^{o}06′13′′.2 (equinox 2000.0), which is 5′′.8 east and 6′′.6 south of the nucleus of UGC 4671. A KAIT image taken on Apr. 8.2 showed nothing at the position of SN 2000dv (limiting mag ∼ 19.5).

SUPERNOVA 2000dk IN NGC 382 K. Krisciunas, University of Washington, reports photometry of SN

2000dk, derived from images taken with the Manastash Ridge Observatory 0.76-m reflector near Ellensburg, WA, using Dophot to obtain raw PSF magnitudes (uncertainties ± 0.03 mag): Sept. 20.45 UT, V = 15.87, B − V = +0.08, V − R = +0.17, V − I = +0.18. From imagery obtained by Krisciunas and R. McMillan with the Apache Point Observatory 3.5-m reflector (and reduced with aperture photometry using IRAF), on Oct. 5.26 they find V = 16.02 ± 0.02, B − V = 0.54 ± 0.02.

SUPERNOVA 2000do IN NGC 6754 Unfiltered CCD magnitude estimates by C. Bembrick, Mt. Tarana Ob-

servatory: Sept. 30.576 UT, 15.0 ± 0.3; Oct. 1.380, 15.1; 2.475, 15.2; 3.375, 15.2; 4.379, 15.2.

2000 October 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green